OpenStudy (anonymous):

how long is leg x on the triangle below 45 one corner 45 the other corner and 90 the other corner and one side is 5' what is x

OpenStudy (anonymous):

you can use the following rule ^_^ :\[sinA/a = sinB/b\] and call your sides , a, b and c. Then solve normally. You have side a and angle A + angle B, then you can find side x= b :)

OpenStudy (anonymous):

First draw the triangle, then call your angles A,B,C and the sides, a, b, c , give it a try now, you'll figure it out ^_^

OpenStudy (anonymous):

i dont now the answer never did this math before can u tell me the answer

OpenStudy (anonymous):

Steps :) : given: A = 45 degrees side a = 5 Angle B = 45 degrees side b = x ( and we want to find x here right? ^_^ ) SinA/a = SinB/b Sin45/5 = Sin45/x then cross multiply and you'll get : \[(\sqrt{2}/2)(x) = (\sqrt{2}/2)(5)\] solve for x now and you'll get : x = 5 Correct me if I'm wrong. Or we could've in other ways concluded that side x = 5 since we have 2 angles that are equal A = B , so the sides are equal too a = b =5 , did you understand the idea? ^_^